# PFT: Phonon Fine-tuning for Machine Learned Interatomic Potentials

### v2

@@ 5 Related Work
Several works have proposed the use of Hessians for training interatomic potentials. Fang et al. (2024) first demonstrated the ability to use AD for phonon prediction through the use of an extended graph, eliminating previously aforementioned issues introduced with self-interaction under finite displacement calculations. They also demonstrate the training of MLIP models on hessians of small organic molecules. Gangan et al. (2025) demonstrate gradient-based optimization of the classical Stillinger-Weber and EDIP potentials to align with DFT-computed phonon and elastic constant calculations. Amin et al. (2025) introduced a method of model distillation by training smaller MLIP on the analytical Hessians of larger, more accurate models. They also randomly sample columns of the Hessian, but use the Jacobian-vector product of direct-force prediction (i.e. non-energy conserving) models for training.
Lastly, (Burger et al., 2025) proposed the direct prediction hessians in small organic molecules, bypassing the need for AD or finite-difference calculations. While this may be beneficial in terms of computational cost, models would not be able to benefit from pre-training on the large quantities of available DFT calculations. Furthermore, tasks such as molecular dynamics, which require lower-order forces, or thermal conductivity which require higher third-order force constants, would be infeasible to conduct.
@@ 6 Discussion
We also highlight practical considerations for training on phonon calculations. Training necessitates access to higher-order derivatives through automatic differentiation, e.g. with JAX. Phonon datasets are biased towards relaxed equilibrium structures, and training on these alone may not resolve errors far from equilibrium, motivating continual learning strategies such as co-training.
@@ Appendix A Appendix
A.1 Energy discrepancy between PBE MDR and MPtrj datasets
@@
A.2 Hyper-parameters
Table A.1 shows the hyper-parameters and method for selection for the Nequix PFT models.
Table A.1: Hyper-parameters used and rationale behind selection
Hyper-parameter Nequix MP PFT Nequix MP PFT (no cotrain) Notes/Rationale
Pretrained Model Nequix MP Nequix MP See Koker et al. (2025) for architecture details.
Learning rate 0.0003 0.0003 Selected from $\{0.003,0.001,0.0003,0.0001\}$ based on validation performance early in training.
Optimizer AdamW AdamW Standard optimizer.
Weight decay 0.001 0.001 From Nequix MP.
Phonon energy weight $\lambda_{E}$ 0 20 From Nequix MP.
Phonon force weight $\lambda_{F}$ 20 20 From Nequix MP.
Phonon stress weight $\lambda_{\sigma}$ 5 5 From Nequix MP.
Phonon force constant weight $\lambda_{\Phi}$ 100 100 Not tuned, selected to be larger than other weights.
Cotrain energy weight $\lambda_{E}$ 500 n/a Started with Nequix MP value $\times$ 10, increased until MPtrj validation energy does not diverge.
Cotrain force weight $\lambda_{F}$ 200 n/a Nequix MP value $\times$ 10.
Cotrain stress weight $\lambda_{\sigma}$ 50 n/a Nequix MP value $\times$ 10.
Cotrain ratio $K$ 4 n/a Trade off between training time and overfitting to phonon data; 4 cotraining steps for every phonon step is reasonable.
Batch size 16 16 Maximum for GPU memory.
Number of epochs 200 200 Not tuned, based on GPU budget. May benefit from longer training, but validation metrics are mostly converged at this duration.
A.3 Elastic properties
We also benchmark the performance effects of PFT on materials elastic properties A.2. For the benchmark we use MatCalc’s Elasticity Calc(Liu et al., 2024) based the settings as described in and the ml-peg package (Kasoar et al., 2025; Batatia et al., 2025b) whereby the elastic tensor and derived properties like the bulk $K$ and shear $G$ modulus are obtained. Voigt-Reuss-Hill (VRH) average is used to report the final values for which we compute the metrics. We observe that the error on shear modulus is lowered for the PFT model while bulk modulus is worsened. Overall, performance on both modulus is improved for the PFT model. We also report the failure rate for the models for which MatCalc calculations fail to converge..

### v3

@@ 5 Related Work
Several works have proposed the use of analytical Hessians with interatomic potentials. Fang et al. (2024) first demonstrated the ability to use AD for phonon prediction through the use of an extended graph, eliminating previously aforementioned issues introduced with self-interaction under finite displacement calculations. They also demonstrate the training of MLIP models on Hessians of small organic molecules. Gangan et al. (2025) demonstrate gradient-based optimization of the classical Stillinger-Weber and EDIP potentials to align with DFT-computed phonon and elastic constant calculations. Amin et al. (2025) introduced a method of model distillation by training smaller MLIP on the analytical Hessians of larger, more accurate models. They also randomly sample columns of the Hessian, but use the Jacobian-vector product of direct-force prediction (i.e. non-energy conserving) models for training.
Lastly, Burger et al. (2025) proposed the direct prediction Hessians in small organic molecules, bypassing the need for AD or finite-difference calculations. While this may be beneficial in terms of computational cost, models would not be able to benefit from pre-training on the large quantities of available DFT calculations. Furthermore, tasks such as molecular dynamics, which require lower-order forces, or thermal conductivity which require higher third-order force constants, would be infeasible to conduct.
@@ 6 Discussion
We also highlight practical considerations for training on phonon calculations. Training necessitates access to higher-order derivatives through automatic differentiation, e.g. with JAX, which may not be accessible for all universal models.
@@ Acknowledgements
Impact Statement
The goal of this work is to advance research in the field of machine learning for computational materials science. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Appendix
A.1 Full Hessian vs. stochastic HVP training
[…]
As the time and space complexity of the full Hessian is $O(N^{2})$ and HVP is $O(N)$ with respect to number of atoms, this performance gap only grows with more atoms. With larger models and radial cutoffs that are typically used for universal MLIPs and larger supercells needed for accurate phonon calculations, the time and memory usage of training on the full analytical Hessian quickly becomes computationally infeasible.
Figure A.1: Full Hessian vs stochastic HVP with Nequix 90K. We train a small 90K parameter model from scratch on the force constants of individual materials, using both the full Hessian and the stochastic HVP. Each column shows the unit cell, the training time and GPU memory consumption on an NVIDIA RTX A5500, and the full Hessian MAE throughout training.
A.2 Energy discrepancy between PBE MDR and MPtrj datasets
@@
A.3 Hyper-parameters
A.3.1 Training the Nequix OAM base model
Following Fu et al. (2025), the Nequix OAM model is trained by first pretraining on OMat24 (Barroso-Luque et al., 2024), and then fine-tuning on a combination of subsampled Alexandria (sAlex) (Schmidt et al., 2024) and eight copies of MPtrj (Deng et al., 2023; Jain et al., 2013). OMat24 and sAlex contain over 100M and 10M DFT calculations respectively. We use the same training setup as the original Nequix MP model (Koker et al., 2025), using the torch backend with OpenEquivariance kernels (Bharadwaj et al., 2025). Changes to hyper-parameters outlined in Table A.1. The OMat pretraining was about 250 A100 hours, and the OAM fine-tuning was an additional 40 A100 hours.
Table A.1: Hyper-parameters used for training OMat24 and OAM base models and rationale behind selection.
Hyper-parameter Nequix OMat Nequix OAM Notes/Rationale
Train data OMat24 MPtrj + sAlex
[…]
Warmup epochs 0.1 0 From Nequix MP; no warmup for fine-tuning.
# of epochs 6 3 Models seem to mostly converge at this duration.
A.3.2 Training PFT models
Table A.2 shows the hyper-parameters and method for selection for the Nequix PFT models, which are trained using the JAX backend on a single A100 GPU.
Table A.2: Hyper-parameters used for PFT models and rationale behind selection.
Hyper- parameter Nequix MP PFT Nequix MP PFT (no co-train) Nequix OAM PFT Notes/Rationale
Base model Nequix MP Nequix MP Nequix OAM See Koker et al. (2025) for architecture details.
Learning rate 0.0001 0.0001 0.0001 Selected from {0.003, 0.001, 0.0003, 0.0001} based on validation performance early in training.
Optimizer AdamW AdamW AdamW Standard optimizer.
Weight decay 0.001 0.001 0.001 From Nequix MP.
PFT $\lambda_{E}$ 0 20 0 From Nequix MP.
PFT $\lambda_{F}$ 20 20 20 From Nequix MP.
PFT $\lambda_{\sigma}$ 5 5 5 From Nequix MP.
PFT $\lambda_{\Phi}$ 100 100 100 Not tuned, selected to be larger than other weights.
Co-train train MPtrj n/a MPtrj + sAlex Same as training data for base model.
Co-train val. MPtrj n/a sAlex Same as validation data for base model.
Co-train $\lambda_{E}$ 500 n/a 750 Started with Nequix MP value $\times$ 10, increased until co-train validation energy does not diverge over time (see Fig. 3).
Co-train $\lambda_{F}$ 200 n/a 200 Nequix MP value $\times$ 10.
Co-train $\lambda_{\sigma}$ 50 n/a 50 Nequix MP value $\times$ 10.
Co-train ratio $K$ 4 n/a 4 Trade off between training time and overfitting to phonon data; 4 cotraining steps for every phonon step is reasonable.
Batch size 16 16 16 Maximum for GPU memory ($\timesA1$00 80GB).
# of epochs 200 200 200 Not tuned, based on GPU budget. May benefit from longer training, but validation metrics are close to converged at this duration.
A.4 Elastic properties
We also benchmark the performance effects of PFT on materials elastic properties, shown in Table A.3. For the benchmark we use the MatCalc Elasticity Calc (Liu et al., 2024) based on the settings as described in the ml-peg package (Kasoar et al., 2025; Batatia et al., 2025b) whereby the elastic tensor and derived properties like the bulk $K$ and shear $G$ modulus are obtained. Voigt-Reuss-Hill (VRH) average is used to report the final values for which we compute the metrics. We observe that the error on shear modulus is lowered for the PFT model while bulk modulus is worsened. Overall, average performance on both modulus is improved for the PFT model. We also report the failure rate for the models for which MatCalc calculations fail to converge.